A recent study found that the average cost of an emergency room visit was $1,200 with a standard deviation of $250. If a patient is selected at random, what is the probability that the person’s visit will cost more than $1,600?


1
Expert's answer
2021-01-19T01:40:15-0500

This task will be solved with assumption of normal distribution.

"μ=1200 \\

σ = 250 \\

z = \frac{x-μ}{σ} \\

= \frac{1600-1200}{250} \\

= 1.60 \\

P(X>1600) = P(Z>1.60) \\

= 1 -P(Z<1.60) \\

= 1 -0.9452 \\

= 0.0548"

Answer: 0.0548 or 5.48 %